OpenStudy (anonymous):

Help please! What are the vertices, foci, and asymptotes of the hyperbola with the equation 16x2 – 4y2 = 64?

OpenStudy (mertsj):

Get it in the right form by dividing each term by 64

OpenStudy (anonymous):

is y=(sqrt)((x-2)(x+2)) right?

OpenStudy (mertsj):

\[\frac{16x^2}{64}-\frac{4x^2}{64}=\frac{64}{64}\]

OpenStudy (tkhunny):

...or \(4y^{2}\)...

OpenStudy (mertsj):

\[\frac{x^2}{4}-\frac{y^2}{16}=1\]

OpenStudy (mertsj):

Does that help?

OpenStudy (anonymous):

Yes! Thanks a bunch! I was completely stuck.

OpenStudy (mertsj):

a=2, b=4

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

So what is the problem?